Could someone help me with this problem.
g(x)=1-sin^(2)X where x is in radians
the function of g can be changed so it can also be described as h(x)=a sin[b(x+(pi/4)]+d
state the values of a, b, c, d.
Thanks for your help.

Answer by stanbon(75887)

g(x)=1-sin^(2)X where x is in radians
the function of g can be changed so it can also be described as h(x)=a sin[b(x+(pi/4)]+d
state the values of a, b, c, d.
--------
The equivalent equation is:
y=(1/2)sin(2(x+pi/4))+(1/2)
a=(1/2), b=2, c is missing but it is pi/4, d=(1/2)
Cheers,
San H.